The reciprocal of two more than a number is three fourths of the reciprocal of the number itself. What is the number?
Answer by checkley77(12844)

1/(x+2)=3/4*1/x
1/(x+2)=3/4x
4x=3(x+2)
4x=3x+6
4x-3x=6
x=6 answer.
Proof:
1/(6+2)=3/4*1/6
1/8=3/24
1/8=1/8